OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. Find w if P = 68 and l = 10.

OpenStudy (anonymous):

2w=p-2l putting valuse of p&w 2w=68-2*10 2w=68-20 2w=48 w=24

OpenStudy (anonymous):

good answer.

OpenStudy (shaik0124):

w=24

OpenStudy (anonymous):

this is all about substitutions,as in,we are going to replace all the known values of the symbols in the equation: if p=2(l+w) p/2=l+w therefore w=(p/2)-l thus;making substitutions we have; 68=2(w+10) 34=10+w w=34-10 w=24